OpenStudy (anonymous):

simplify difference (4w^2-3w-7)-(6w^2+2w-3)

OpenStudy (anonymous):

remove parenthesis then combine like terms

OpenStudy (mathstudent55):

Remove the first set of parentheses because they are unnecessary. To remove the second set of parentheses, you must change the sign of every term in the second set of parentheses since there is a negative sign outside. Then combine like terms.

OpenStudy (anonymous):

10w^2+5w+4 ?? this is the answer I got

OpenStudy (mathstudent55):

(4w^2-3w-7)-(6w^2+2w-3) = 4w^2 - 3w - 7 - 6w^2 - 2w + 3 = -2w^2 - 5w - 4

OpenStudy (mathstudent55):

(4w^2-3w-7)-(6w^2+2w-3) = 4w^2 - 3w - 7 - 6w^2 - 2w + 3 = ^ ^ ^ |______|____|_______ These terms stay the same bec the first set of parentheses was unnecessary.

OpenStudy (mathstudent55):

(4w^2-3w-7)-(6w^2+2w-3) = 4w^2 - 3w - 7 - 6w^2 - 2w + 3 = ^ |_________ 6w^2 inside parentheses became -6w^2

OpenStudy (mathstudent55):

(4w^2-3w-7)-(6w^2+2w-3) = 4w^2 - 3w - 7 - 6w^2 - 2w + 3 = ^ |________ 2w inside parenthses became -2w

OpenStudy (mathstudent55):

(4w^2-3w-7)-(6w^2+2w-3) = 4w^2 - 3w - 7 - 6w^2 - 2w + 3 = ^ |______________ -3 inside parentheses became +3

OpenStudy (anonymous):

ok thank you

OpenStudy (mathstudent55):

wlcm